# Judging the Judges: Evaluating Alignment and Vulnerabilities in LLMs-as-Judges

### v2

@@ Abstract
Offering a promising solution to the scalability challenges associated with human evaluation, the LLM-as-a-judge paradigm is rapidly gaining traction as an approach to evaluating large language models (LLMs). However, there are still many open questions about the strengths and weaknesses of this paradigm, and what potential biases it may hold. In this paper, we present a comprehensive study of the performance of various LLMs acting as judges. [^1] Source code is available at https://github.com/UMass-Meta-LLM-Eval/llm_eval.git We leverage TriviaQA as a benchmark for assessing objective knowledge reasoning of LLMs and evaluate them alongside human annotations which we found to have a high inter-annotator agreement. Our study includes nine judge models and nine exam-taker models – both base and instruction-tuned. We assess the judge models’ alignment across different model sizes, families, and judge prompts. Among other results, our research rediscovers the importance of using Cohen’s kappa as a metric of alignment as opposed to simple percent agreement, showing that judges with high percent agreement can still assign vastly different scores. We find that both Llama-3 70B and GPT-4 Turbo have an excellent alignment with humans, but in terms of ranking exam taker models, they are outperformed by both JudgeLM-7B and the lexical matching method Contains, which have up to 34 points lower human alignment. Through error analysis and various other studies, including the effects of instruction length and leniency bias, we hope to provide useful lessons for using LLMs as judges in the future.
@@ 1 Introduction
Figure 1: (a) Scores assigned to all exam-taker models by the various judge models. (b) Average percent agreement (blue line) and Kappa scores (red bars) of judge models with human judges. Error bars annotate standard deviation across exam-taker models. Alignment is poor for most judge models, but both Llama3 70B and GPT-4 Turbo have Cohen’s Kappa coefficient that are indicative of excellent alignment (79 and 84, respectively), though still well below the human alignment score of 96.
@@
To evaluate LLMs, various methods have been proposed, that typically fall into one of two broad categories. Benchmarks such as MMLU (Hendrycks et al., 2021), TruthfulQA (Lin et al., 2021), and GSM8K (Cobbe et al., 2021) are used to evaluate specific capabilities of LLMs in an automated manner. Additionally, leaderboards like Chatbot Arena (Chiang et al., 2024) and Open LLM Leaderboard (Beeching et al., 2023) assign ranks to models considering pair-wise rankings of LLM outputs, done by humans or, in some cases, automated evaluation methods. Since both strategies involve evaluating free-form text responses generated by the LLMs, even in the first case, evaluating the responses is often just as challenging as generating them (see e.g. Chang et al., 2023).
@@
One proposed solution to this problem is to use multiple-choice question (MCQ) benchmarks such as MMLU, and compare the log-probabilities of the potential answers rather than evaluating the generated answer directly. However, the MCQ paradigm ly limits the range of abilities that can be evaluated, and the setup increasingly diverges from how LLMs are used in practice. Alternatively, the use of lexical matching methods such as exact match (EM) or n-gram overlap to evaluate the responses are practical and cost-efficient approaches, but are susceptible to false negatives and often fail to adequately distinguish between responses with subtle differences that change their semantic meaning. This issue is exacerbated when evaluating instruction-tuned “chat” models that are fine-tuned to carry out conversations with humans in natural language, since their responses tend to be more verbose (Saito et al., 2023; Renze and Guven, 2024). For these reasons, human evaluation remains the gold standard for evaluating LLM responses. However, human evaluation is expensive, time-consuming, and often impractical in many use cases. As a result, it has increasingly become common practice to evaluate LLM responses using another LLM as a judge model (Lin et al., 2021; Islam et al., 2023; Chiang and Lee, 2023; Liusie et al., 2024). While there are promises of alignment between LLM judges and humans (Sottana et al., 2023; Zheng et al., 2024), there are also many open questions about the strengths and weaknesses of the paradigm.
@@
In this work, we study the properties of LLMs as judges, comparing them with humans and automated evaluation methods. Contrary to prior work, we focus on a relatively ‘clean’ scenario in which inter-human agreement is as high as 96%, allowing us to distinguish ambiguity and subjectivity in the task itself from potential issues with the judge models. Using the knowledge benchmark TriviaQA (Joshi et al., 2017) as our playground, we investigate how nine different judge models with varying architectures and sizes judge nine different exam-taker models. Our main findings are:
@@
  • Even in relatively straightforward setups, only the best models are suitable as judges. Out of the nine judge models we considered, only GPT-4 Turbo and Llama-3 70B showed very high alignment with humans, though even those judges’ alignment is still well behind the human alignment coefficient for the task (Figure 1).
@@
  • Cohen’s kappa distinguishes judges better than percent alignment. In some cases, high percent agreement can still give very divergent scores (Figure 2).
@@
  • Even Cohen’s kappa is not all telling. While GPT-4 Turbo and Llama-3 both have alignment scores that are considered excellent for discriminating different exam-taker models, their results are comparable to alternative cheaper approaches such as JudgeLM-7B and contains which have much lower alignment scores but more systematic biases (Figure 3)
@@
Through detailed error analysis, we uncover additional insights into judge performance. Improved alignment appears to be driven by improved recall rates and reduced false negatives. However, judge models struggle with under-specified answers and tend to be lenient, affecting their evaluation consistency. They are also sensitive to the length and quality of prompts. And, surprisingly, even when the judge models are asked to evaluate an answer matching verbatim with a reference answer, many judge models still sometimes fail to evaluate it correctly.
@@ 2 Related work
Various recent studies have used or considered using LLMs as judges for tasks such as evaluating story generation (Chiang and Lee, 2023), retrieval-augmented generation (Es et al., 2023), visual QA (Mañas et al., 2024), code comprehension (Zhiqiang et al., 2023), multilingual evaluation (Hada et al., 2023) and more general open-ended tasks (Zheng et al., 2024). Zhang et al. (2024) and Sottana et al. (2023) propose ways to standardise LLM evaluations and the role that judge models might play in such solutions. Several studies have demonstrated that state-of-the-art LLMs such as GPT-4 Turbo exhibit high alignment with human judgments (Sottana et al., 2023; Zheng et al., 2024), though others also illustrate that the paradigm is not yet without faults. Zeng et al. (2023) propose a benchmark for evaluating the performance of LLMs as judges, and other approaches have been proposed to improve LLM judges such that they are aligned well with humans (Shankar et al., 2024; Zhu et al., 2023).
Despite promising results in various settings, judge models still suffer from known issues of current LLMs such as hallucinations and factual errors (Ye et al., 2023; Turpin et al., 2023) and difficulty in following complex instructions (Li et al., 2023b; He et al., 2024). Furthermore, various studies have reported challenges such as position bias (Pezeshkpour and Hruschka, 2023; Zheng et al., 2023; Wang et al., 2023), verbosity bias (Saito et al., 2023) in their preferences, confusing evaluation criteria (Hu et al., 2024), or focusing more on the style and grammar compared to factuality (Wu and Aji, 2023). Recently, Liusie et al. (2024) have shown that LLMs perform better in comparative assessment compared to absolute scoring, which can be used for reliably measuring the relative performance of models (Liu et al., 2024) and creating classifiers for pairwise grading (Huang et al., 2024).
@@
Judge models (instruction-tuned) Llama-2 (7B, 13B, and 70B), Llama-3 (8B and 70B),
@@
Table 1: The exam-taker models and judge models we use in our experiments. We consider a wide variety of judge models; to get a comprehensive overview of their (potential) biases, we consider exam-taker models of various sizes and types.
@@ 3 Methodology
To evaluate the strengths and weaknesses of the LLM-as-a-judge paradigm, we focus on a relatively controlled setup, in which judge models assess answers of exam-taker models on the knowledge benchmark TriviaQA (Joshi et al., 2017). We consider nine judge models and nine exam-taker models. More details about the dataset and models are provided in Appendix B.
Benchmark
As our testbed, we use the TriviaQA dataset (Joshi et al., 2017), consisting of 95K question-answer pairs sourced from 14 trivia and quiz league websites. Each question in the train and validation set is annotated with a list of short answers containing a minimal set of facts and evidence documents collected from Wikipedia and the Web. For our experiments, we use the validation set of the unfiltered partition of the benchmark, using the short answers as reference answers. We use the training set for few-shot examples. Since experiments require manual annotation of the exam-taker model responses, we use a random sample of $400$ questions from the dataset.
@@
To understand the strengths and weaknesses of different judges, we benchmark pre-trained (base) and instruction-tuned (chat) exam-taker models across a wide variety of model sizes and examine the quality of the evaluations from different judge models. In particular, we consider Llama-2 (Touvron et al., 2023) in 7B, 13B, and 70B parameter sizes for both base and chat versions, Mistral 7B (Jiang et al., 2023) base and chat versions, and GPT-4 Turbo [^2] Accessed via the OpenAI API between Mar 19th, 2024 and May 19th, 2024. (Achiam et al., 2023) as the exam-taker models. The prompts for the exam-taker models contain five few-shot examples of (question, answer) pairs from the TriviaQA training set. The prompts for the instruction-tuned models additionally include a command signaling the model to answer the given question in a succinct manner similar to the provided examples. The prompts are provided in Appendix C.
@@
To get a comprehensive view of the strengths and weaknesses of judge models across different model sizes and architectures, we use instruction-tuned versions of Llama-2 (Touvron et al., 2023) in 7B, 13B, and 70B sizes, Llama-3 (AI@Meta, 2024) in 8B and 70B sizes, Mistral 7B (Jiang et al., 2023), GPT-4 Turbo (Achiam et al., 2023), Gemma 2B (Gemma Team et al., 2024), and JudgeLM 7B (Zhu et al., 2023) as judges. The judges are instructed to respond with only a single word, ‘‘correct’’ or ‘‘incorrect’’. The prompts can be found in Appendix D. An overview of al exam-taker models and judge models is shown in Table 1. For ease of reading, the judge models are depicted in a different font than the exam-taker models.
@@
We use two metrics to quantify alignment between judges: percent agreement and Cohen’s kappa coefficient (Cohen, 1960). Percent agreement expresses a simple percentage of the samples on which two annotators agree. Cohen’s kappa, denoted as $\mathbf{\kappa}$, is an alignment metric that also takes into account the possibility of chance agreement. It is generally considered to provide a more robust measure of alignment. Details about the computation of both metrics are given in Appendix E.
@@
As a ground-truth assessment, we obtain human annotations for each exam-taker model answer. The inter-human alignment is calculated between three human judges using the answers to 600 questions by Llama-2 7B base, randomly sampled from the TriviaQA dataset (Joshi et al., 2017); the human guidelines can be found in Appendix F. We then determine collective “Human Judgment” through a majority vote. We found that the average alignment among human evaluators with the majority vote had a Cohen’s kappa score [^3] The values of the Cohen’s kappa coefficient have been scaled by 100×100$\times$100 × for easier comparison with percent agreement of $96.36\pm 1.67$, and the average percent agreement was $98.33\%\pm 0.76\%$. Given this near-perfect alignment score, we consider only one human evaluator per sample for the rest of our experiments, to reduce the overall cost of human annotations. The set of questions for which we obtain human annotations is identical for each exam-taker model.
@@ 4 Results
Figure 2: Delta evaluation score is calculated by taking judge score difference with human judgment; y-axes are in log scale. Fig (a) shows a skewed distribution for percent agreement and delta evaluation score. In Fig (b) we observe that highly aligned LLM judges with kappa > 0.8 exhibit low variability in scores. Conversely, judges with kappa < 0.8 demonstrate more variability, impacting their reliability.
In this section we discuss our main results, primarily focusing on the relationship between evaluations by various judge models and human evaluations (§ 4.1), and how that impacts their usability (§ 4.2). To do so, we evaluate their alignment with human judgment and assess how differently they rank the nine exam-taker models compared to humans. In Section 5, we further analyse their precision and recall to further investigate the types of errors that can be made by various judge models. Details about compute requirements and others costs for experiments are given in Appendix G.
@@ 4.1 Alignment between judge models and humans
We first compute $\kappa$ scores and percent agreement between the evaluations of each judge model and the human annotators. Figure 1 shows that while alignment is poor for most judge models, both Llama-3 70B and GPT-4 Turbo have $\mathbf{\kappa}$ scores – 79 and 84, respectively – that are considered to indicate excellent alignment. Nevertheless, there still is a significant disparity between human judgment and judge models: GPT-4 Turbo is 12 points behind human judgment. Notably, contains has a higher $\mathbf{\kappa}$ score than half of the judge models, while EM has the lowest alignment among all judges.
Cohen’s kappa vs percent agreement
@@
In most cases, we observe that Cohen’s kappa and percent agreement are following the same trend, with the exception of the values for Gemma 2B and EM. Gemma 2B shows higher percent agreement compared to EM, yet it yields the lowest $\mathbf{\kappa}$ score within the ensemble. Furthermore, there is a significant difference in the actual values. For the percent agreement of judge models, we note a 30-point difference between human judgment and EM, while Cohen’s kappa exhibits a more substantial 53-point gap. This is also visible in the general decline of alignment scores: while Llama-3 8B has a $\kappa$ score of only 62, its percent agreement is still well above 80%. Overall, $\mathbf{\kappa}$ scores offer a more precise representation of diminishing trends in judge models compared to percent agreement.
Alignment vs assigned score
@@
In Figure 2, we show the variation in scores assigned by the judge models to various exam-taker models for different values of percent agreement (Figure 2(a)) and Cohen’s kappa (Figure 2(b)). We can see that for $\mathbf{\kappa}$ > 80, the evaluation scores by judge models are close to the human evaluation scores for most of the judges, with a difference of up to 5 points in their assigned scores (complete results table provided in Appendix H). For percent agreement we observe deviations of up to 20 points in the evaluation scores for similar kappa alignment. Furthermore, we observe that the deviation from human-judgements can be quite distinct for different exam-taker models. In Figure 1(a), Gemma 2B, for instance, sometimes assigns higher scores than humans, and sometimes much lower. In the next section, we further explore this particular pattern.
4.2 Exploring systematic patterns in judge models

### v3

@@ Abstract
Offering a promising solution to the scalability challenges associated with human evaluation, the LLM-as-a-judge paradigm is rapidly gaining traction as an approach to evaluating large language models (LLMs). However, there are still many open questions about the strengths and weaknesses of this paradigm, and what potential biases it may hold. In this paper, we present a comprehensive study of the performance of various LLMs acting as judges [^1] Source code is available at https://github.com/UMass-Meta-LLM-Eval/llm_eval Investigating thirteen judge models of different model sizes and families, judging answers of nine different ‘exam-taker models’ – both base and instruction-tuned – we find that only the best (and largest) models achieve reasonable alignment with humans. However, they are still quite far behind inter-human agreement and their assigned scores may still differ with up to 5 points from human-assigned scores. In terms of their ranking of the nine exam-taker models, instead, also smaller models and even the lexical metric contains may provide a reasonable signal. Through error analysis and other studies, we identify vulnerabilities in judge models, such as their sensitivity to prompt complexity and length, and a tendency toward leniency. The fact that even the best judges differ from humans in this comparatively simple setup suggest that caution may be wise when using judges in more complex setups. Lastly, our research rediscovers the importance of using alignment metrics beyond simple percent alignment, showing that judges with high percent agreement can still assign vastly different scores.
@@ 1 Introduction
Figure 1: Average scores assigned by judge models and alignment with human judges. (a) Scores assigned to all exam-taker models by the various judge models. (b) Average percent agreement (blue line) and Scott’s pi scores (red bars) of judge models with human judges (black line). Error bars annotate standard deviation across exam-taker models. Llama3 70B, Llama3.1 70B and GPT-4 Turbo have Scott’s pi coefficient that are indicative of excellent alignment, but are still well below the human alignment score.
@@
To evaluate LLMs, various methods have been proposed, typically falling into one of two broad categories. First, benchmarks such as MMLU (Hendrycks et al., 2021), TruthfulQA (Lin et al., 2021), or GSM8K (Cobbe et al., 2021) are used to evaluate specific capabilities of LLMs in an automated manner. Additionally, leaderboards like Chatbot Arena (Chiang et al., 2024) and Open LLM Leaderboard (Beeching et al., 2023) assign ranks to models considering pair-wise rankings of LLM outputs, done by humans or, in some cases, automated evaluation methods. Since both strategies involve evaluating free-form text responses generated by the LLMs, even in the first case, evaluating the responses is often just as challenging as generating them (see e.g. Chang et al., 2023; Bavaresco et al., 2024).
@@
One proposed solution to this problem is to use multiple-choice question (MCQ) benchmarks such as MMLU, and compare the log-probabilities of the potential answers rather than evaluating the generated answer directly. However, the MCQ paradigm limits the range of abilities that can be evaluated, and the setup increasingly diverges from how LLMs are used in practice. Alternatively, the use of lexical matching methods such as exact match (EM) or n-gram overlap to evaluate the responses are practical and cost-efficient approaches, but are susceptible to false negatives and often fail to adequately distinguish between responses with subtle differences that change their semantic meaning. This issue is exacerbated when evaluating instruction-tuned “chat” models that are fine-tuned to carry out conversations with humans in natural language, since their responses tend to be more verbose (Saito et al., 2023; Renze & Guven, 2024). For these reasons, human evaluation remains the gold standard for evaluating LLM responses.
@@
However, human evaluation is expensive, time-consuming, and often impractical in many use cases. As a result, it has increasingly become common practice to evaluate LLM responses using another LLM as a judge model (Lin et al., 2021; Islam et al., 2023; Chiang & Lee, 2023; Liusie et al., 2024). While there are promises of alignment between LLM judges and humans (Sottana et al., 2023; Zheng et al., 2024), there are also many open questions about the strengths and weaknesses of the paradigm. In this work, we study the properties of LLMs as judges, comparing them with humans and automated evaluation methods. Contrary to prior work, we focus on a clean scenario in which human alignment is very high, allowing us to distinguish ambiguity and subjectivity in the task itself from potential issues with the judge models. Using the knowledge benchmark TriviaQA (Joshi et al., 2017) as our playground, we investigate how thirteen different judge models with varying architectures and sizes judge nine different exam-taker models. Our main findings are:
@@
  • Even in clean and straightforward setups, only the best models have high alignment scores. Out of the thirteen judge models we considered, only GPT-4 Turbo, Llama-3.1 70B and Llama-3 70B showed very high alignment with humans. Also for those judges, though, alignment is still well behind the human alignment coefficient for the task (Figure 1).
@@
  • Scott’s $\mathbf{\pi}$ distinguishes judges better than percent alignment. In terms of percent alignment, judges are rarely discriminable, while Scott’s $\mathbf{\pi}$ provides a more informative signal. In some cases, high percent agreement can still give scores that differ 10-20 points from the human-assigned scores (Figure 2).
@@
  • Also Scott’s $\mathbf{\pi}$ is not all telling. While GPT-4 Turbo and Llama-3 both have alignment scores that are considered excellent, their scores still differ up to 5 points from human-assigned scores. Furthermore, when it comes to discriminating different exam-taker models, their results are comparable to alternative cheaper approaches such as Mistral 7B and contains, which have much lower alignment scores but more consistent biases (Figure 3).
@@
Through detailed analysis (§ 5), we uncover additional insights into judge performance. Improved alignment appears to be driven by improved recall rates and reduced false negatives. However, judge models struggle with under-specified answers and tend to be lenient, affecting their evaluation consistency. They are also sensitive to the length and quality of prompts. And, surprisingly, even when the judge models are asked to evaluate an answer matching verbatim with a reference answer, many judge models still sometimes fail to evaluate it correctly.
@@ 2 Related work
Various recent studies have used or considered using LLMs as judges for tasks such as evaluating story generation (Chiang & Lee, 2023), retrieval-augmented generation (Es et al., 2023), visual QA (Mañas et al., 2024), code comprehension (Zhiqiang et al., 2023), multilingual evaluation (Hada et al., 2023) and more general open-ended tasks (Zheng et al., 2024). Zhang et al. (2024) and Sottana et al. (2023) propose ways to standardise LLM evaluations and the role that judge models might play in such solutions. Several studies have demonstrated that state-of-the-art LLMs such as GPT-4 Turbo exhibit high alignment with human judgments (Sottana et al., 2023; Zheng et al., 2024), though others also illustrate that the paradigm is not yet without faults. Zeng et al. (2023) propose a benchmark for evaluating the performance of LLMs as judges, and other approaches have been proposed to improve LLM judges such that they are aligned well with humans (Shankar et al., 2024; Zhu et al., 2023).
Despite promising results in various settings, judge models still suffer from known issues of current LLMs such as hallucinations and factual errors (Ye et al., 2023; Turpin et al., 2023) and difficulty in following complex instructions (Li et al., 2023b; He et al., 2024). Furthermore, various studies have reported challenges such as position bias (Pezeshkpour & Hruschka, 2023; Zheng et al., 2023; Wang et al., 2023), verbosity bias (Saito et al., 2023) in their preferences, confusing evaluation criteria (Hu et al., 2024), or focusing more on the style and grammar compared to factuality (Wu & Aji, 2023). Recently, Liusie et al. (2024) have shown that LLMs perform better in comparative assessment compared to absolute scoring, which can be used for reliably measuring the relative performance of models (Liu et al., 2024) and creating classifiers for pairwise grading (Huang et al., 2024).
@@
Table 1: Exam-taker models and judge models We consider a wide variety of exam-taker models and judge models; to get a in-depth overview of their abilities, we consider exam-taker models of various sizes & types.
@@
Judge models (instruction-tuned) Llama-2 (7B, 13B, 70B),
Llama-3 (8B, 70B), Llama-3.1 (8B, 70B),
@@
Judge models (lexical) Exact Match (EM), Contains
@@ 3 Methodology
To evaluate the strengths and weaknesses of the LLM-as-a-judge paradigm, we focus on a comparatively controlled setup, in which judge models assess answers of exam-taker models on the knowledge benchmark TriviaQA (Joshi et al., 2017). With this methodological design, it is possible to focus on the abilities of the judges in isolation, without having to address human disagreement and error at the same time. In this section, we elaborate the main aspects of our methodology.
Evaluation data
As our testbed, we use the TriviaQA dataset (Joshi et al., 2017), consisting of 95K question-answer pairs sourced from 14 trivia and quiz league websites. Each question in the train and validation set is annotated with a list of short answers containing a minimal set of facts and evidence documents collected from Wikipedia and the Web. For our experiments, we use the validation set of the unfiltered partition of the benchmark, using the short answers as reference answers. We use the training set for few-shot examples. Since experiments require manual annotation of the exam-taker model responses, we use a random sample of $400$ questions from the dataset. In Appendix I, we show with a bootstrapping test that this sample size has low variance for our main result. Through experiments described in § 3, we establish that humans have high agreement on judgements of answers given to the questions in the benchmark.
@@
To understand the strengths and weaknesses of different judges, we consider answers of pre-trained (base) and instruction-tuned (chat) ‘exam-taker models’ across a wide variety of model sizes. In particular, we consider Llama-2 (Touvron et al., 2023) in 7B, 13B, and 70B parameter sizes for both base and chat versions, Mistral 7B (Jiang et al., 2023) base and chat versions, and GPT-4 Turbo [^2] Accessed via the OpenAI API between Mar 19th, 2024 and Sep 20, 2024. (Achiam et al., 2023) as the exam-taker models. The prompts for the exam-taker models contain five few-shot examples of (question, answer) pairs from the TriviaQA training set. The prompts for the instruction-tuned models additionally include a command signaling the model to answer the given question in a succinct manner similar to the provided examples. The prompts are provided in Appendix D.
@@
To get a comprehensive view of the strengths and weaknesses of judge models across different model sizes and architectures, we use instruction-tuned versions of Llama-2 (Touvron et al., 2023) in 7B, 13B, and 70B sizes, Llama-3 (AI@Meta, 2024) in 8B and 70B sizes, Llama-3.1 (Dubey et al., 2024) in 8B and 70B sizes, Mistral 7B (Jiang et al., 2023), GPT-4 Turbo (Achiam et al., 2023), Gemma 2B (Gemma Team et al., 2024), and JudgeLM 7B (Zhu et al., 2023) as judges. To maintain parity with human and judge evaluation, judge prompts were built from human guidelines in Appendix G. The judges are instructed to respond with only a single word, ‘‘correct’’ or ‘‘incorrect’’. An overview of all exam-taker models and judge models is shown in Table 1. For ease of reading, the judge models are depicted in a different font than the exam-taker models.
@@
We use two metrics to quantify alignment between judges: percent agreement and Scott’s Pi coefficient (Scott, 1955). [^3] In an earlier version of this paper, we used Cohen’s kappa (Cohen, 1960) to measure alignment. It has since come to our attention that – despite it’s widespread use – this metric has some well-documented theoretical issues (e.g. Pontius & Millones, 2011; Chicco et al., 2021). For the interested reader, we elaborate on these issues in Appendix B. Percent agreement expresses a simple percentage of the samples on which two annotators agree. Scott’s Pi, denoted as Scott’s $\mathbf{\pi}$, is an alignment metric that corrects for chance agreement between two annotators and is considered to provide a more robust measure of alignment. Details about the computation of both metrics are given in Appendix F.
@@
As a ground-truth assessment, we obtain human annotations for each exam-taker model answer. The inter-human alignment is calculated between three human judges using the answers to 1200 randomly sampled questions answers; the human guidelines can be found in Appendix G. We then determine collective “Human Judgment” through a majority vote. The average alignment among human evaluators with the majority vote had a Scott’s $\mathbf{\pi}$ of $96.2\pm 1.07$,444coefficient scaled by $100$ for easier comparison with percent alignment. and the average percent agreement was $98.52\%\pm 0.42\%$. The details of this experiment are mentioned in Appendix A. Given this near-perfect alignment score, we consider only one human evaluator per sample for the rest of our experiments, to reduce the overall cost of human annotations. The set of questions for which we obtain human annotations is identical for each exam-taker model.
@@ 4 Results
Figure 2: Difference with human evaluation scores versus alignment metric. The delta evaluation score is the difference between the judge and the human score; y-axes are in log scale. Percent alignment (left) shows a very skewwed distribution, making it difficult to distinguish models. Scott’s $\mathbf{\pi}$ (left) provides a clearer difference between models, and is more indicative of deviation of the gold score.
In this section we discuss our main results, primarily focusing on the relationship between evaluations by various judge models and human evaluations (§ 4.1), and how that impacts their usability (§ 4.2). To do so, we evaluate their alignment with human judgment and assess how differently they rank the nine exam-taker models compared to humans. In Section 5, we further analyse their precision and recall to further investigate the types of errors that can be made by various judge models. Details about compute requirements and others costs for experiments are given in Appendix H.
@@ 4.1 Alignment between judge models and humans
We start by computing Scott’s $\mathbf{\pi}$ scores and percent agreement between the evaluations of each judge model and the human annotators. We show the result in Figure 1. We observe that percent alignment is high for virtually all models, with the exception of Gemma 2B and EM. Scott’s $\mathbf{\pi}$, on the other hand, has low values for most models, though its value is in the high 80s for Llama-3 70B, Llama-3.1 70B and GPT-4 Turbo. Nevertheless, there still is a significant disparity between human judgment and judge models: the best scoring judge, Llama-3 70B, is 8 points behind human judgment. Notably, EM has the most variance in alignment, while Gemma 2B has the lowest alignment amongst all judges.
@@
In most cases, we observe that Scott’s $\mathbf{\pi}$ and percent agreement are following the same trend, with the exception of the values for Gemma 2B and EM. Gemma 2B shows higher percent agreement compared to EM, yet it yields the lowest Scott’s $\mathbf{\pi}$ score within the ensemble. For the percent agreement of judge models, we note a 26-point difference between human judgment and EM, while Scott’s $\mathbf{\pi}$ exhibits a more substantial 64-point gap. This is also visible in the general decline of alignment scores: while Llama-3 8B has a Scott’s $\mathbf{\pi}$ score of only 59, its percent agreement is still well above 80%. Overall, Scott’s $\mathbf{\pi}$ appears to be better able of discriminating various judge models, showing more divergence across the tested judges.
@@
To understand how indicative the two alignment metrics are of the expected accuracy of the overall judgement of the models, we plot, for each judge model and exam-taker model, the difference between the score assigned by the judge and the score assigned by a human. In the figure, we can see that for Scott’s $\mathbf{\pi}$ values higher than 80, the evaluation scores are comparatively close to the human evaluation scores, with a difference of up to 5 points in their assigned scores (complete results table provided in Appendix J). For percent alignment, on the other hand, even judges that have more than 90% may still differ more than 10 points in their assigned score. Interestingly, the deviation from human-judgements for a single judge model can be quite different depending on the exam-taker model. In Figure 1(a), Gemma 2B, for instance, sometimes assigns higher scores than humans, and sometimes much lower. In the next section, we further explore this particular pattern.
4.2 Exploring consistent patterns in judge models